1. A motorcycle begins driving from at rest position. It accelerates at 8 km/hr2 for 20 seconds, then travels at a steady speed for another 20 seconds, all in the same direction. How much distance has it covered since it started driving?


2. An object is dropped from a cliff, the time it takes to hit the ground is 3 seconds. What is the height of the cliff? What is the velocity of the object before it hits the ground?


1
Expert's answer
2022-06-01T12:49:18-0400

1.

"v=v_0+at=0+6.17*10^{-4}*20=0.012\:\rm m/s"

"d=d_1+d_2=at_1^2/2+vt_2\\
=6.17*10^{-4}*20^2/2+0.012*20=0.37\:\rm m"

2.

"h=gt^2/2=9.81*3^2/2=44\:\rm m"

"v=gt=9.81*3=29\:\rm m/s"